# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v2

@@ D.1 Silver nanoparticles at SwissFEL
D.2 Two-color experiment at the European XFEL
The experiment was performed at the Nano-sized Quantum Systems (NQS) endstation of the Small Quantum Systems (SQS) instrument [73, 74, 75], located on the SASE3 branch of the European XFEL [76]. The XFEL was operated at a repetition rate of $10\text{\,}\mathrm{Hz}$ with an electron energy of $16\text{\,}\mathrm{GeV}$.
@@ D.2 Two-color experiment at the European XFEL
[…]
Appendix E Retrieving the shape of helium droplets
Figure 7: Reconstruction of a xenon-doped helium nanodroplet from a two-color diffraction pattern. The figure presents the same data shown in Fig. 4a–c. However, in this case, the two reconstructed frames in panels b) and c) are not rescaled to match a common size bar and retain their original pixel dimensions.
@@ Appendix E Retrieving the shape of helium droplets
For a spherical particle, the scattered field can be calculated exactly and analytically using Mie theory, which provides solutions to Maxwell’s equations [77]. The scattering signal can be fully computed with knowledge of the droplet size $D$, the radiation wavelength $\lambda$, the refractive index of the material $n^{He}(\lambda)$ at the given photon energy, and the energy density of the light field.
For a diffraction signal from the same pure-helium droplet, composed of the incoherent sum of two different colors, Ref. [35] shows that it is possible to retrieve the size of the droplet and the intensity of the two pulses by fitting the recorded radial intensity profile as the incoherent sum of two Mie-calculated intensities, one for each color. In this procedure, the droplet size $D$ is assumed constant between the two scattering events, while the wavelengths of the two pulses, $\lambda^{A}=1.25\,\unit\nano$ and $\lambda^{B}=1.04\,\unit\nano$, as well as the corresponding refractive indices for helium, $n^{He}(\lambda^{A})$ and $n^{He}(\lambda^{B})$ [49], are known a priori.
@@
The physical system is thus compatible with a compact spherical helium nanodroplet for both the pump and the probe pulses, as supported by both experimental observations and theoretical considerations. Furthermore, it is possible to extract the droplet size in a precise and consistent way by fitting Mie solutions to Maxwell’s equations for a signal produced by two different photon energies. For a thorough investigation of the physical system and the fitting procedure, please refer to Ref. [35]. This provides the a-priori information necessary for the use of the DCDI method during the imaging process with Dichography.
@@
As discussed in Sec. A.3, the iterative approach to Dichography retains the real-space operations of the conventional iterative phase retrieval algorithms for CDI. The DCDI method is, in practice, an additional real-space constraint and can be directly implemented in the algorithm workflow (see Fig. 5). The only required additional step is to properly adapt the droplet size of the density profile to the corresponding spatial resolution for the two frames, $\rho^{A}$ and $\rho^{B}$.
@@
In fact, the spatial extension of a single pixel in the reconstruction corresponds to the half-period resolution [7]. In the small-angle approximation [6], the spatial extension of a single reconstruction pixel $\Delta$ can be calculated as:
$\Delta=\frac{\lambda}{2\sin(\theta_{\text{max}})}$ (7)
@@
where $\lambda$ is the radiation wavelength and $\theta_{\text{max}}$ is the angle measured from the center of the detector to its side edge. For a two-color reconstruction with Dichography, the two frames correspond to two views of the same sample obtained with different colors. As such, the corresponding pixel size $\Delta$ differs. In particular, the values for the $1.0\text{\,}\mathrm{keV}$ ($\lambda=1.25\,\unit\nano$) pulse and the $1.2\text{\,}\mathrm{keV}$ ($\lambda=1.04\,\unit\nano$) pulse can be calculated via Eq. (7) as $\Delta^{\text{$1.0\text{\,}\mathrm{keV}$}}=5.57\,\unit\nano$ and $\Delta^{\text{$1.2\text{\,}\mathrm{keV}$}}=4.64\,\unit\nano$.
@@
As a consequence, when reconstructing a sample of size $D$, the $1.0\text{\,}\mathrm{keV}$ reconstruction will have an extent of $\nicefrac{{D}}{{\Delta^{\text{$1.0\text{\,}\mathrm{keV}$}}}}$ pixels, while the $1.2\text{\,}\mathrm{keV}$ reconstruction will extend over $\nicefrac{{D}}{{\Delta^{\text{$1.2\text{\,}\mathrm{keV}$}}}}$ pixels. To better highlight this aspect, we report in Fig. 7 the same two-color reconstruction shown in Fig. 4a-c, where the two retrieved frames are not scaled to match the spatial extension; i.e., they are cropped from the reconstructed matrices $\rho^{A}$ and $\rho^{B}$ with the same pixel dimensions. On the one hand, this representation helps the reader visualize the lower resolution of the $1.0\text{\,}\mathrm{keV}$ pulse. On the other hand, it highlights that the appearance of two similar structures in the two frames cannot derive from a ghosting effect. In fact, the arising of a reconstruction feature in the wrong frame wouldn’t match the proper spatial resolution, since the two structures differ in pixel extent.

### v3

@@ V.4.1 Silver nanoparticles at SwissFEL
V.4.2 Two-color experiment at the European XFEL
The experiment was performed at the Nano-sized Quantum Systems (NQS) endstation of the Small Quantum Systems (SQS) instrument [81, 82, 83], located on the SASE3 branch of the European XFEL [84] (proposal 2857). The XFEL was operated at a repetition rate of $10\text{\,}\mathrm{Hz}$ with an electron energy of $16\text{\,}\mathrm{GeV}$.
@@ V.4.2 Two-color experiment at the European XFEL
V.5 Retrieving the shape of helium droplets
@@
For a spherical particle, the scattered field can be calculated exactly and analytically using Mie theory, which provides solutions to Maxwell’s equations [85]. The scattering signal can be fully computed with knowledge of the droplet size $D$, the radiation wavelength $\lambda$, the refractive index of the material $n^{\text{He}}(\lambda)$ at the given photon energy, and the energy density of the light field.
For a diffraction signal from the same pure-helium droplet, composed of the incoherent sum of two different colors, Ref. [38] shows that it is possible to retrieve the size of the droplet and the intensity of the two pulses by fitting the recorded radial intensity profile as the incoherent sum of two Mie-calculated intensities, one for each color. In this procedure, the droplet size $D$ is assumed constant between the two scattering events, while the wavelengths of the two pulses, $\lambda^{A}=$1.25\text{\,}\mathrm{nm}$\$ and $\lambda^{B}=$1.04\text{\,}\mathrm{nm}$\$, as well as the corresponding refractive indices for helium, $n^{\text{He}}(\lambda^{A})$ and $n^{\text{He}}(\lambda^{B})$ [54], are known a priori.
@@
The physical system is thus compatible with a compact spherical helium nanodroplet for both the pump and the probe pulses, as supported by both experimental observations and theoretical considerations. Furthermore, it is possible to extract the droplet size in a precise and consistent way by fitting Mie solutions to Maxwell’s equations for a signal produced by two different photon energies. For a thorough investigation of the physical system and the fitting procedure, please refer to Ref. [38]. This provides the a priori information necessary for the use of the DCDI method during the imaging process with Dichography.
@@
As discussed in Sec. V.1.3, the iterative approach to Dichography retains the real-space operations of the conventional iterative phase retrieval algorithms for CDI. The DCDI method is, in practice, an additional real-space constraint and can be directly implemented in the algorithm workflow (see Fig. 5). The only required additional step is to properly adapt the droplet size of the density profile to the corresponding spatial resolution for the two frames, $\rho^{A}$ and $\rho^{B}$.
V.6 Representation of two-color reconstructions
Figure 7: Two-color reconstructions of the xenon-doped helium nanodroplets shown in Fig. 2, here reported with the original sampling. The figure presents the same data as shown in Fig. 2. Each two-frame reconstruction is reported twice. On the left, the densities are reported with enhanced contrast, similarly to Fig. 2, to highlight the inner xenon structures. On the right, the original scattering density is reported. In this second case, the color scale is shared between the two frames to also report the imbalance between the scattering contributions.
@@ V.6 Representation of two-color reconstructions
The spatial extension of a single pixel in the reconstruction corresponds to the half-period resolution [7]. In the small-angle approximation [6], the spatial extension of a single reconstruction pixel $\Delta$ can be calculated as:
$\Delta=\frac{\lambda}{2\theta_{\text{max}}}$ (13)
@@
where $\lambda$ is the radiation wavelength and $\theta_{\text{max}}$ is the angle measured from the center of the detector to its side edge. For a two-color reconstruction with Dichography, the two frames correspond to two views of the same sample obtained with different colors. As such, the corresponding pixel size $\Delta$, i.e., the sampling in real space, differs. In particular, the values for the $1.0\text{\,}\mathrm{keV}$ ($\lambda^{A}=$1.25\text{\,}\mathrm{nm}$\$) pulse and the $1.2\text{\,}\mathrm{keV}$ ($\lambda^{B}=$1.04\text{\,}\mathrm{nm}$\$) pulse can be calculated via Eq. (13) as $\Delta^{\text{$1.0\text{\,}\mathrm{keV}$}}=$5.57\text{\,}\mathrm{nm}$\$ and $\Delta^{\text{$1.2\text{\,}\mathrm{keV}$}}=$4.64\text{\,}\mathrm{nm}$\$.
@@
As a consequence, when reconstructing a sample of size $D$, the $1.0\text{\,}\mathrm{keV}$ reconstruction will have an extent of $\nicefrac{{D}}{{\Delta^{\text{$1.0\text{\,}\mathrm{keV}$}}}}$ pixels, while the $1.2\text{\,}\mathrm{keV}$ reconstruction will extend over $\nicefrac{{D}}{{\Delta^{\text{$1.2\text{\,}\mathrm{keV}$}}}}$ pixels. To better highlight this aspect, we report in Fig. 7 the same two-color reconstructions as shown in Fig. 2, where the two retrieved frames are not scaled to match the spatial extension; i.e., they are cropped from the reconstructed matrices $\rho^{A}$ and $\rho^{B}$ with the same pixel dimensions. On the one hand, this representation helps the reader visualize the lower resolution of the $1.0\text{\,}\mathrm{keV}$ pulse. On the other hand, it highlights that the appearance of two similar structures in the two frames cannot derive from a ghosting effect. In fact, the arising of a reconstruction feature in the wrong frame wouldn’t match the proper spatial resolution, since the two structures differ in pixel extent. In addition, Fig. 7 reports the two scattering densities without the enhancement of the contrast for the xenon structures and with a common color scale between the two frames. The reconstructed densities are reported with the correct scaling of the helium droplet $\rho_{\bigcirc}$ and the xenon doping $\rho_{\chi}$, i.e., $\rho=\rho_{\bigcirc}+\rho_{\chi}$ . This allows the reader to better appreciate the relatively low contrast of the xenon filaments within the droplets, and thus better evaluate the noise visible in the reconstructions shown in enhanced contrast in Fig. 2. Furthermore, it is worth noting that the scattering density $\rho$, outcome of the reconstruction, is proportional to the actual electronic density $\rho_{e}$ and the field amplitude $\sqrt{I}$, i.e., $\rho\propto\sqrt{I}\cdot\rho_{e}$. The visibly fainter $1.2\text{\,}\mathrm{keV}$ reconstruction in Fig. 7a comes indeed from the lower intensity contribution of the $1.2\text{\,}\mathrm{keV}$ pulse in the scattering, which corresponds to a weaker $1.2\text{\,}\mathrm{keV}$ field in the specific position of the imaged droplet within the two focal regions.